OpenStudy (anonymous):

How do I find out what the ratio of 4/5 to 7/15 is?

OpenStudy (anonymous):

(7/15)/(4/5) = (7/15)*(5/4) = (7/3)*(1/4) = ?

OpenStudy (anonymous):

I know the answer is 1 5/7 but not sure on they came about getting this answer